OpenStudy (anonymous):

Suppose that 19 inches of wire costs 57 cents. At the same rate, how many inches of wire can be bought for 36 cents?

OpenStudy (kl0723):

you're given the conversion factor here 19 inches of wire = 57 cents, put it on paper and solve the problem using this

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

On the chart it says 1 cm = .3937 in but I'm not sure how it would help here

OpenStudy (kl0723):

\[36 cents \times \frac{ 19 inches }{ 57 cents }\]

OpenStudy (anonymous):

I got 12 as my answer

OpenStudy (kl0723):

that is it :)

OpenStudy (anonymous):

Another question

OpenStudy (kl0723):

36 cents on top cancel with 36 cents on the bottom and you're left with units of (inches of wire)

OpenStudy (kl0723):

I meant 57 cents on the bottom :)

OpenStudy (kl0723):

only the units... not the numbers

OpenStudy (anonymous):

So would I have to do the same step with this: Greg drove 605 miles in 11 hours. At the same rate, how many miles would he drive in 13 hours?

OpenStudy (anonymous):

605 x 11 over 13?

OpenStudy (kl0723):

yes

OpenStudy (kl0723):

but look... the numbers don't matter... is about the units, you have to make sure that you cancel the units you don't need and stay with units u need for your problem

OpenStudy (anonymous):

605 cancels itself? So just divide the 11 and 13 then?

OpenStudy (kl0723):

the numbers do not cancel... the units do :)

OpenStudy (anonymous):

I got 511.923 :/

OpenStudy (kl0723):

\[13 h \times \frac{ 605mi }{ 11 h } = ?\]

OpenStudy (anonymous):

Oh >.>

OpenStudy (kl0723):

set up like this... h = hours... so h on top cancels with h on the bottom and you're left with units of mi = miles :)

OpenStudy (kl0723):

and now do the operation 13 x 605... all divided by 11

OpenStudy (anonymous):

Does it always have to be h/h? It was 715

OpenStudy (kl0723):

hours divided by hours cancels the units... if you're trying to know how many miles... then make sure you're left with (mi) units by canceling hours

OpenStudy (anonymous):

It makes sense now, thank you

OpenStudy (kl0723):

good :)